OpenStudy (anonymous):

A parallelogram has sides measuring 23.8 m and 35.3 m. The height corresponding to the 23.8-m base is 9.9 m. Find the height, to the nearest tenth of a meter, corresponding to the 35.3-m base. A. 8317.4 m B. 14.7 m C. 6.7 m D. none of these

OpenStudy (anonymous):

I know the answers not B

OpenStudy (anonymous):

use proportions \[\frac{ 23.8 }{ 9.9 }=\frac{ 35.3 }{ x }\]

OpenStudy (anonymous):

u know how to do that?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

well look at my equation that i put up there and multiple 35.3 and 9.9

OpenStudy (anonymous):

then divide by 23.8

OpenStudy (anonymous):

it comes out to be 14.684 meters

OpenStudy (anonymous):

yeah i did that and i got 14.6836134. which rounds to 14.7 but im retaking the exam and i missed this question the first time and i put b and they said it was wrong

OpenStudy (anonymous):

well then the question is wrong, but your right :)

OpenStudy (anonymous):

ok ill email my teacher and have him look at it. the assignment was computer graded so its probably just a mistake. thank you

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

That is wrong.

OpenStudy (anonymous):

the area of parallelogram is : side times height corresponding to that side, so 23,8*9,9=35,3*h h=6.6764 Correct answer is C. 6.7